OpenStudy (anonymous):

Find the equation of the circle of radius sqr(10) that are tangent to 3x+y=6 at (3,-3)

OpenStudy (anonymous):

Okay, so we have a point on the surface of the circle, right? That's (3,-3). And we have its radius. The equation of this circle is (x-h)^2+(y-k)^2=10, where (h,k) is the center and \[\sqrt{10}\] is the radius.

OpenStudy (anonymous):

So, what we really need to find is the center.

OpenStudy (anonymous):

working it out, I was able to get k=1/3h-4

OpenStudy (anonymous):

then plug that into the original eq of (3-h)^2+(-3-k)^2=10

OpenStudy (anonymous):

Since we know that the line represented by 3x+y=6 is tangent to the circle, then the line perpendicular to 3x+y=6 at (3,-3) will pass right through the center of the circle. We also know that it will do so at a point that is \[\sqrt{10}\] distant from (3,-3).

OpenStudy (anonymous):

Donutlicious, that's only going to get you something like 1=1 or some such.

OpenStudy (anonymous):

So, to get a line perpendicular to 3x+y=6, let's put it into point-slope format: y=-3x-6 Alright, so a line perpendicular to that one would be y = 3x, since it would have the opposite slope. However, we want the line to intersect at the point (3,-3), so we'll need to add some constant b and solve for it: y = 3x + b -3 = 3(3) + b -3 - 9 = b b = -12

OpenStudy (anonymous):

@jonnybot the perpendicular line would have slope \[\frac{1}{3}\] but i think you are on the right track

OpenStudy (anonymous):

Satellite can u come help me when u r finished? ;D

OpenStudy (anonymous):

Okay, so we've got out perpendicular line! Now we know that the center of the circle is sqrt(10) away from (3,-3) and that it lies on the line y=3x-12. Using the distance formula between the center of the circle and the line, we know that: sqrt(10) = sqrt{(3-h)^2 + (-3-k)^2} Where (h,k) is the center of the circle).

OpenStudy (anonymous):

Ah I see. I'm jotting this all down as we go through this. Thank you for your time by the way :)

OpenStudy (anonymous):

Also, I'm pretty sure satellite is wrong, and that the perpendicular line would have slope of 3.

OpenStudy (anonymous):

hey don't put down satellite!

OpenStudy (anonymous):

pretty sure the perpendicular line has slope \[\frac{1}{3}\] and equation \[y+3=\frac{1}{3}(x-3)\] or \[y=\frac{1}{3}x-4\] so any point on the line will look like \[(x. \frac{1}{3}x-4)\] and if it is \[\sqrt{10}\] away from \[(3,-3)\] then you know that \[(x-3)^2+\frac{1}{3}(x-1)^2=10\] solve to get \[x=0, x=\frac{24}{5}\]

OpenStudy (anonymous):

@jonnybot line perpendicular to a line with slope m has slope \[-\frac{1}{m}\] and so if your line has slope -3, perpendicular line will have slope \[\frac{1}{3}\]

OpenStudy (anonymous):

And Satellite... look at http://www.wolframalpha.com/input/?i=graph+y%3D-3x+and+y%3D3x . Tell me those lines aren't perpendicular. Then look at http://www.wolframalpha.com/input/?i=graph+y%3D-3x+and+y%3D%281%2F3%29x and try to tell me that they are.

OpenStudy (anonymous):

By the way, I'm not trying to talk smack here or anything.

OpenStudy (anonymous):

Wait... satellite is right... http://www.wolframalpha.com/input/?i=graph+y%3D-3x+and+y%3D%281%2F3%29x+from+x%3D-1+to+x+%3D1+and+y+%3D+-1+to+y+%3D+1 I was deluded. :)

OpenStudy (anonymous):

Ah man, he left :( I still needed help.